# Error Correction Code Transformer: From Non-Unified to Unified

### v3

@@ 5.3 Computational Complexity Analysis
Table 2: Comparison of trainable parameters, MACs, training time, inference time, and mask density among ECCT, FECCT, and UECCT for various codes, all with identical parameters $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=2048$
@@
Compared to the approach in  [24, 25], while the proposed method excels in decoding versatility and computational complexity, it may fall short in terms of memory efficiency, power consumption, and computational resource utilization when compared to non-learning solutions. This inefficiency could limit its potential for deployment. To address these issues, future implementations could leverage techniques such as parameter sharing, low-rank factorization, and quantization, aiming to reduce complexity and enhance efficiency  [45, 46].
@@ 6 Conclusions
We introduced a unified, code-agnostic Transformer decoding architecture, designed to integrate various linear block codes seamlessly and offer superior decoding performance. Our approach leverages the self-attention mechanism inherent in Transformer models, enabling effective handling of sequential data and capturing long-range dependencies critical for decoding tasks. Through rigorous analysis and extensive experimental validation, our results indicate that the unified Transformer decoding architecture can significantly advance the state-of-the-art in decoding performance, potentially setting new benchmarks for future research and practical applications.
In conclusion, the unified Transformer decoding architecture presented in this paper signifies a significant step forward in the field of error correction coding. It paves the way for the development of more adaptable, efficient, and robust decoding solutions, poised to meet the stringent requirements of 6G and beyond. As the field progresses, we anticipate that the continued integration of AI technologies will unlock new possibilities, driving innovation in wireless communication systems. Future work can aim to further reduce hardware implementation complexity and enhance computational efficiency using techniques such as quantization and compression of neural network models.

### v4

@@ 5.3 Computational Complexity Analysis
Table 2: Comparison of trainable parameters, MACs, training time, inference time, and mask density among ECCT, FECCT, and UECCT for various codes, all with identical parameters $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=2048$.
@@
Compared to the approach in [22, 23], while the proposed method excels in decoding versatility and computational complexity, it may fall short in terms of memory efficiency, power consumption, and computational resource utilization when compared to non-learning solutions. This inefficiency could limit its potential for deployment. To address these issues, future implementations could leverage techniques such as parameter sharing, low-rank factorization, and quantization, aiming to reduce complexity and enhance efficiency [44, 45, 46].
@@ 6 Conclusions
We proposed a unified Transformer-based decoder that seamlessly integrates multiple linear block codes within a single framework. By introducing standardized unit, a unified attention module, and a sparse mask leveraging the parity-check matrix’s sparsity, we achieved enhanced decoding accuracy and reduced computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$. This work delivers a high-performance, low-complexity solution, addressing hardware overhead and scalability challenges in next-generation wireless systems like 6G. Future work will focus on improving memory efficiency and power consumption through techniques such as pruning and quantization to enhance deployability in resource-constrained environments.